# Testing identifying assumptions in Tobit Models

### v2

@@ 9. Conclusion
In this paper, we develop sharp testable equalities for the classic Tobit and IV-Tobit models that can detect all observable violations of the model’s assumptions. The results are shown to extend to many other popular Tobit-type “two-part” models. By converting these sharp equalities into conditional moment inequalities, we propose a testing procedure that detects violations of the Tobit model assumptions on a grid on the joint support of the outcome (and treatment) variables leveraging inference results from Chernozhukov, Lee, and Rosen (2013) and the implementation from Chernozhukov et al. (2015).
Simulation results suggest the test performs well for reasonably sized samples (larger than 5000 observations). The test is conservative for smaller samples, overejecting the null hypothesis of model validity. Simulations indicate that the test is powerful to detect violations of the exogeneity assumption for the treatment/instrument that affect the point estimates and inference. Finally, the proposed test exhibits good performance for violations of the distributional assumptions about the error structure.
We provide a brief review of existing models that could be implemented under weak/alternative assumptions when the Tobit model is rejected. Furthermore, we propose a simple model that partially identifies the parameter of interest by relying solely on linear index and monotone treatment selection restrictions, a standard assumption from the partial identification literature (Manski and Pepper, 2000).
We illustrate our methods on data from Lee (1995). We replicate qualitatively the results in the original paper and the proposed test for validity of the IV-Tobit model rejects the null hypothesis in this empirical application. We estimate our proposed lower bound for the effect of household income from sources other than the wife’s labor on their labor supply, which does not rely on the normality of latent errors or treatment exogeneity. While we can rule out that an extra 1,000 dollars in other household income reduce female labor supply by more than 4.2 hours per year, we cannot rule out that the effect is zero.
@@ Appendix A Sharpness for the classic Tobit case
Let $\alpha_{1},\alpha_{0}$ be identified. Suppose 3.1- 3.3 hold. Let $f_{\tilde{U}|D}=\frac{1}{\sqrt{2\pi}}e^{-1/2\tilde{U}^{2}}$. Let $\tilde{Y}^{*}=\alpha_{0}+\alpha_{1}D+\tilde{U}$ and $\tilde{Y}=\max(0,\tilde{Y}^{*})$. From $f_{\tilde{U}|D}$ we can see assumption 1 holds, since $f_{\tilde{U}|D}=f_{\tilde{U}}$. Also note $f_{\tilde{U}}$ is the $N(0,1)$ density, thus assumption 2 holds. Considering the positive values for $\tilde{Y}$, for any constants $c_{0},c_{1}$ such that $0<c_{0}<c_{1}$, then,
@@
$=$ $P(c_{0}\leq\alpha_{0}+\alpha_{1}d+\tilde{U}\leq c_{1}|D=d)$
$=$ $P(c_{0}-\alpha_{0}-\alpha_{1}d\leq\tilde{U}\leq c_{1}-\alpha_{0}-\alpha_{1}d|D=d)$
$=$ $P(c_{0}-\alpha_{0}-\alpha_{1}d\leq\tilde{U}\leq c_{1}-\alpha_{0}-\alpha_{1}d)$
$=$ $\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)-\Phi(c_{0}-\alpha_{0}-\alpha_{1}d)$
@@
The first three equalities follow from the definitions of $\tilde{Y}^{*}$ and $\tilde{Y}$. The fourth and fifth steps use the specific choice for the probability density of $\tilde{U}$, which implies independence from $D$ and normality respectively. The last step uses the relationship between the observed data and the model, established in 3.1. Similar reasoning applies for $P(c_{1}\leq\tilde{Y}|D=d)=P(c_{1}\leq\alpha_{0}+\alpha_{1}d+\tilde{U}|D=d)=1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)=P(c_{1}\leq Y|D=d)$.
@@
$P(\tilde{Y}=0|D=d)$ $=$ $P(\tilde{Y}^{*}\leq 0|D=d)=P(\alpha_{0}+\alpha_{1}d+\tilde{U}\leq 0|D=d)$
$=$ $P(\tilde{U}\leq-\alpha_{0}-\alpha_{1}d|D=d)=P(\tilde{U}\leq-\alpha_{0}-\alpha_{1}d)$
$=$ $\Phi(-\alpha_{0}-\alpha_{1}d)=1-\Phi(\alpha_{0}+\alpha_{1}d)=P(Y=0|D=d)$
The last step uses the equality established in 3.3. Thus, we characterized the distribution of $\tilde{Y}|D$, which is equal to the one of $Y|D$, with $D$ given. Thus, we can induce the observed distribution $Y,D$.
@@ Appendix B Details of the IV-Tobit case
$=$ $P(c_{0}\leq\beta_{0}+\beta_{1}z+W\leq c_{1},d_{0}\leq\gamma_{0}+\gamma_{1}z+V\leq d_{1},Y^{*}\geq 0|Z=z)$
$=$ $P(c_{0}-\beta_{0}-\beta_{1}z\leq W\leq c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z\leq V\leq d_{1}-\gamma_{0}-\gamma_{1}z,Y^{*}\geq 0|Z=z)$
$=$ $P(c_{0}-\beta_{0}-\beta_{1}z\leq W\leq c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z\leq V\leq d_{1}-\gamma_{0}-\gamma_{1}z)$
@@
B.2. Derivation of equation 3.10
By a similar approach to the derivation of 3.4:
@@
$=$ $P(W\leq-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}\leq V\leq d_{1}-\gamma_{0}-\gamma_{1}z)$
$=$ $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
-- $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
@@ B.3. Proof of Sharpness
Let $\beta_{0},\beta_{1},\gamma_{0},\gamma_{1},\rho$ be identified and equalities 3.4-3.12 hold. Define the joint density of $\left(\tilde{W},\tilde{V}\right)$ conditional on $Z$ as
@@
Note that $f_{\left(\tilde{W},\tilde{V}|Z\right)}=f_{\left(\tilde{W},\tilde{V}\right)}$, thus assumption 3 holds. Similarly, $\left(\tilde{W},\tilde{V}\right)$ follow a bivariate normal distribution as $\begin{pmatrix}\tilde{W}\\ \tilde{V}\end{pmatrix}\sim\mathcal{N}(\mu,\Sigma)$, where $\mu=\begin{pmatrix}0\\ 0\end{pmatrix}$, and $\Sigma=\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}$. Let $\tilde{U}=\tilde{W}-\alpha_{1}\tilde{V}$, which implies $\tilde{U},\tilde{V}$ satisfies assumption 4, given the scale-location normalizations.
@@
The first through third equalities follow from the definitions of $\tilde{Y},\tilde{Y}^{*}$ and $\tilde{D}$. The fourth and fifth steps are consequences of the particular choice for joint density for $\left(\tilde{W},\tilde{V}\right)$ conditional on $Z$. The final equality is given by the relationship between observable data and the latent model structure in Equation 3.4. Similar derivations hold for 3.5-3.9.
@@
The steps of the proof are similar to previous cases and the last equality is given by the relationship between observable data and the latent model structure in Equation 3.10. Similar derivations hold for equalities 3.11-3.12.
@@ Appendix C Type 2 Tobit testable implications
In this section we derive the results of the Type 2 tobit model also known as selection models or Heckman selection type of models (Heckman, 1979). The basic setup with no covariates (which can be extended in several directions and with different distributional assumptions as well as to incorporate treatment endogeneity, as we pointed out for the classic Tobit model in remark 3) is:
@@
Note that $Y$ is missing at $S=0$. Thus, fully characterizing the distribution implies observing the behaviour at the missing point and beyond it.
@@
Where the first and the second equalities are due to the structure of the model described in equation C.8. The third step is due to Assumption 7 and the last one follows from the normalization for normal random variables and Assumption 8.
@@
$=$ $\Phi_{U}(c_{1}-\alpha_{0}-\alpha_{1}d)-\Phi_{U}(c_{0}-\alpha_{0}-\alpha_{1}d)$
-- $\Phi_{U,V}(c_{1}-\alpha_{0}-\alpha_{1}d,-\gamma_{0}-\gamma_{1}z,\rho_{UV})$
$+$ $\Phi_{U,V}(c_{0}-\alpha_{0}-\alpha_{1}d,-\gamma_{0}-\gamma_{1}z,\rho_{UV})$
The first equality follows from the law of total probability. The second through fifth equalities follow from the model’s structure described in Equation C.8, namely $P(Y>0,Y^{*}<0|D=d)=0$ and $Y^{*}=Y,\text{not missing}$ for $S=1$, the structure of $S$, and the latent linear model. The sixth step is due to Assumption 7. The last equality uses Assumption 8 as well as properties of normal random variables. Thus, we can construct a test similarly to the one in the main text, using these equalities in addition to the ones of the form $P(Y\geq c_{2}|D=d,Z=z)$.
Appendix D Results from Remark 3
In this appendix we derive the testable implications of the extensions to the classic Tobit and IV-Tobit discussed in Remark 3. We discuss testable implications assuming the identification of the relevant parameters or distributions.
D.1. Barros et al. (2018) and Carson and Sun (2007)
Barros et al. (2018) proposes a variant of the tobit model with elliptically contoured distributions and a non-zero threshold. At the same time Carson and Sun (2007) proposes a Tobit model with a non-zero threshold. In this section we combine both types of results and report the testable implications with a generic non-zero threshold and a generic known or identifiable parametric distribution function. In addition, extend the latent structure to be also known or an identifiable function up to a vector of parameters but invertible in $U$. In this context, let:
@@ Appendix D Results from Remark 3
D.2. Wooldridge (2005); Honore, Kyriazidou, and Powell (2000); Honoré (1993)
Consider the following dynamic version of the Tobit model, which is related to Wooldridge (2005); Honore, Kyriazidou, and Powell (2000); Honoré (1993). Here, we specify the conditional behaviour of $c_{i}$ in the spirit of Wooldridge (2005) and others.
@@ D.2. Wooldridge (2005); Honore, Kyriazidou, and Powell (2000); Honoré (1993)
[…]
@@
[…]
Which can then be used to construct a test in a similar way as presented in the main text by adding the type of equalities $P(Y_{i,t}\geq c_{2}|D_{i,t}=d,Y_{i,t-1}=y)$, with the caveat that the right-hand side requires numerical integration or an approximation by an estimator when no closed form is available. Such estimator should ensure that the left-hand side of the equality converges at root-$N$ in order for the estimation step of the null model to be asymptotically negligible (Acerenza, Bartalotti, and Kédagni, 2023, Appendix B).
@@ E.1. Non-Learnability for the classic Tobit model
As mentioned in Remark 2, the standard Tobit model is non-verifiable, that is, we cannot learn if the maintained Tobit model is the true data generating process based on the sharp equalities proposed in Section 3.1. One can show that by finding an alternative model that is compatible with the equalities (3.1)-(3.3) in all cases in which the Tobit model could not be disregarded. Suppose that,
$\left\{\begin{array}[]{lcl}\bar{Y}&=&\max(0,\bar{Y}^{*})\\ \\ \bar{Y}^{*}&=&G(\alpha_{0}+\alpha_{1}D+\bar{U})\end{array}\right.$ (E.4)
Where $G(.)$ is some continuous and strictly increasing function such that $G^{-1}(0)=0$ and $\bar{U}$ is a random variable that is $N(0,1)$.
For $c_{0}\leq Y\leq c_{1}$ for a $c_{1},c_{0}>0$ and for any value of the treatment variables $d\in\mathcal{D}$, assume that the equalities (3.1)-(3.3) hold:
@@
Now define $\bar{Y}_{d}=G(\alpha_{0}+\alpha_{1}d+\bar{U})$ and define the following distribution function:
$P(\bar{Y}_{d}\leq c_{1},\bar{U}\leq u|D=d)=\Phi(\min\{(G^{-1}(c_{1})-\alpha_{0}-\alpha_{1}d),u\})$
In the above proposed DGP, the latent model is not linear. However, the DGP is compatible with the data since,
$P(\bar{Y}_{d}\leq 0)$ $=$ $P(\tilde{U}\leq-\alpha_{0}-\alpha_{1}d)=\Phi(-\alpha_{0}-\alpha_{1}d)$ (E.5)
$=$ $1-\Phi(\alpha_{0}+\alpha_{1}d)=P(Y=0|D=d)$
The last equality is due to the testable equalities holding. Similarly, for any $c_{1}>0$, we can get that
$P(G(c_{1})<\bar{Y}_{d})$ $=$ $1-\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)=P(c_{1}<Y|D=d).$ (E.6)
And,
$P(G(c_{0})\leq\bar{Y}_{d}\leq G(c_{1}))$ $=$ $\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)-\Phi(c_{0}-\alpha_{0}-\alpha_{1}d)$ (E.7)
@@
Hence, when the sharp equalities hold, there exists an alternative observationally equivalent model, to the classic tobit that can induce the observed data distribution. Indeed, the observed distribution of $Y$ and the non refuted marginal distribution of $U$, do not imply that we can learn the true distribution of $Y_{1},Y_{0},U$.

### v3

@@ 9. Conclusion
In this paper, we develop sharp testable equalities for the classic Tobit and IV-Tobit models that can detect all observable violations of the model’s assumptions. The results are shown to extend to many other popular Tobit-type “two-part” models. By converting these sharp equalities into conditional moment inequalities, we propose a testing procedure that detects violations of the Tobit model assumptions on a grid on the joint support of the outcome (and treatment) variables, leveraging inference results from chernozhukov_intersection_2013 and the implementation from chernozhukov_implementing_2015.
Simulation results suggest the test performs well for reasonably sized samples (larger than 5000 observations). The test is conservative for smaller samples, over-rejecting the null hypothesis of model validity. Simulations indicate that the test is powerful to detect violations of the exogeneity assumption for the treatment/instrument that affect the point estimates and inference. Finally, the proposed test exhibits good performance for violations of the distributional assumptions about the error structure.
We provide a brief review of existing models that could be implemented under weak/alternative assumptions when the Tobit model is rejected. Furthermore, we propose a simple model that partially identifies the parameter of interest by relying solely on linear index and monotone treatment selection restrictions, a standard assumption from the partial identification literature (manski2000amonotone).
We illustrate our methods on data from lee1995semi. We replicate qualitatively the results in the original paper and the proposed test for validity of the IV-Tobit model rejects the null hypothesis in this empirical application. We estimate our proposed lower bound for the effect of household income from sources other than the wife’s labor on their labor supply, which does not rely on the normality of latent errors or treatment exogeneity. While we can rule out that an extra 1,000 dollars in other household income reduce female labor supply by more than 4.2 hours per year, we cannot rule out that the effect is zero.
@@ Appendix A Sharpness for the classic Tobit case
[…]
@@
$=$ $P(c_{0}\leq\tilde{\alpha}_{0}+\tilde{\alpha}_{1}d+\tilde{U}\leq c_{1}|D=d)$
$=$ $P(c_{0}-\tilde{\alpha}_{0}-\tilde{\alpha}_{1}d\leq\tilde{U}\leq c_{1}-\tilde{\alpha}_{0}-\tilde{\alpha}_{1}d|D=d)$
$=$ $P\left(\frac{c_{0}}{\sigma}-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}d\leq\frac{\tilde{U}}{\sigma}\leq\frac{c_{1}}{\sigma}-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}d\right)$
$=$ $\Phi\left(\frac{c_{1}}{\sigma}-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}d\right)-\Phi\left(\frac{c_{0}}{\sigma}-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}d\right)$
@@
The first three equalities follow from the definitions of $\tilde{Y}^{*}$ and $\tilde{Y}$. The fourth and fifth steps use the specific choice for the probability density of $\tilde{U}$, which implies independence from $D$ and normality respectively. The last step uses the relationship between the observed data and the model, established in Equation (4.1).
@@
$P(\tilde{Y}=0|D=d)$ $=$ $P(\tilde{Y}^{*}\leq 0|D=d)=P(\tilde{\alpha}_{0}+\tilde{\alpha}_{1}d+\tilde{U}\leq 0|D=d)$
$=$ $P(\tilde{U}\leq-\tilde{\alpha}_{0}-\tilde{\alpha}_{1}d|D=d)=P(\tilde{U}\leq-\tilde{\alpha}_{0}-\tilde{\alpha}_{1}d)$
$=$ $\Phi\left(-\frac{\tilde{\alpha}_{0}}{\sigma}-\frac{\tilde{\alpha}_{1}}{\sigma}d\right)=1-\Phi\left(\frac{\tilde{\alpha}_{0}}{\sigma}+\frac{\tilde{\alpha}_{1}}{\sigma}d\right)=P(Y=0|D=d)$
The last step uses the equality established in Equation 4.2. Thus, we characterized the distribution of $\tilde{Y}|D$, which is equal to the one of $Y|D$, with $D$ given. Thus, we can induce the observed distribution $Y,D$.
@@ Appendix B Details of the IV-Tobit case
$=$ $P(c_{0}\leq\tilde{\beta}_{0}+\tilde{\beta}_{1}z+\tilde{W}\leq c_{1},d_{0}\leq\tilde{\gamma}_{0}+\tilde{\gamma}_{1}z+\tilde{V}\leq d_{1},Y^{*}\geq 0|Z=z)$
$=$ $P(c_{0}-\tilde{\beta}_{0}-\tilde{\beta}_{1}z\leq\tilde{W}\leq c_{1}-\tilde{\beta}_{0}-\tilde{\beta}_{1}z,d_{0}-\tilde{\gamma}_{0}-\tilde{\gamma}_{1}z\leq\tilde{V}\leq d_{1}-\tilde{\gamma}_{0}-\tilde{\gamma}_{1}z,Y^{*}\geq 0|Z=z)$
$=$ $P(c_{0}-\tilde{\beta}_{0}-\tilde{\beta}_{1}z\leq\tilde{W}\leq c_{1}-\tilde{\beta}_{0}-\tilde{\beta}_{1}z,d_{0}-\tilde{\gamma}_{0}-\tilde{\gamma}_{1}z\leq\tilde{V}\leq d_{1}-\tilde{\gamma}_{0}-\tilde{\gamma}_{1}z)$
$=$ $\Phi_{W,V}\left(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma}_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma}_{1}}{\sigma_{\tilde{V}}}z;\rho\right)-\Phi_{W,V}\left(\frac{c_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma}_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma}_{1}}{\sigma_{\tilde{V}}}z;\rho\right)$
- $\Phi_{W,V}\left(\frac{c_{1}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma}_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma}_{1}}{\sigma_{\tilde{V}}}z;\rho\right)+\Phi_{W,V}\left(\frac{c_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma}_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma}_{1}}{\sigma_{\tilde{V}}}z;\rho\right)$
@@
B.2. Derivation of equation (4.9)
By a similar approach to the derivation of (4.3):
@@
$=$ $P(\tilde{W}\leq-\tilde{\beta}_{0}-\tilde{\beta}_{1}z,d_{0}-\tilde{\gamma}_{0}-\tilde{\gamma}_{1}\leq\tilde{V}\leq d_{1}-\tilde{\gamma}_{0}-\tilde{\gamma}_{1}z)$
$=$ $\Phi_{W,V}\left(-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{1}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma}_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma}_{1}}{\sigma_{\tilde{V}}}z;\rho\right)$
- $\Phi_{W,V}\left(-\frac{\tilde{\beta}_{0}}{\sigma_{\tilde{W}}}-\frac{\tilde{\beta}_{1}}{\sigma_{\tilde{W}}}z,\frac{d_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma}_{0}}{\sigma_{\tilde{V}}}-\frac{\tilde{\gamma}_{1}}{\sigma_{\tilde{V}}}z;\rho\right)$
@@ B.3. Proof of Sharpness
[…]
@@
[…]
@@
[…]
@@
The steps of the proof are similar to previous cases and the last equality is given by the relationship between observable data and the latent model structure in Equation 4.9. Similar derivations hold for equalities (4.10)-(4.11).
@@ Appendix C Type 2 Tobit testable implications
In this section, we derive the results of the Type 2 tobit model, also known as selection models or Heckman selection-type models (heckman1979sample). The basic setup with no covariates (which can be extended in several directions and with different distributional assumptions as well as to incorporate treatment endogeneity, as we pointed out for the classic Tobit model in proposition 2) is:
@@
Note that $Y$ is missing at $S=0$. Thus, fully characterizing the distribution implies observing the behavior at the missing point and beyond it.
@@
Where the first and the second equalities are due to the structure of the model described in equation (C.8). The third step is due to Assumption 7 and the last one follows from the normalization for normal random variables and Assumption 8.
@@
$=$ $\Phi_{U}\left(\frac{c_{1}-\alpha_{0}-\alpha_{1}d}{\sigma_{U}}\right)-\Phi_{U}\left(\frac{c_{0}-\alpha_{0}-\alpha_{1}d}{\sigma_{U}}\right)$
- $\Phi_{U,V}\left(\frac{c_{1}-\alpha_{0}-\alpha_{1}d}{\sigma_{U}},\frac{-\gamma_{0}-\gamma_{1}z}{\sigma_{V}},\rho_{UV}\right)$
$+$ $\Phi_{U,V}\left(\frac{c_{0}-\alpha_{0}-\alpha_{1}d}{\sigma_{U}},\frac{-\gamma_{0}-\gamma_{1}z}{\sigma_{V}},\rho_{UV}\right)$
The first equality follows from the law of total probability. The second through fifth equalities follow from the model’s structure described in Equation (C.8), namely $P(Y>0,Y^{*}<0|D=d)=0$ and $Y^{*}=Y,\text{not missing}$ for $S=1$, the structure of $S$, and the latent linear model. The sixth step is due to Assumption 7. The last equality uses Assumption 8 as well as properties of normal random variables. Thus, we can construct a test similarly to the one in the main text, using these equalities in addition to the ones of the form $P(Y\geq c_{2}|D=d,Z=z)$.
Appendix D Results from Proposition 2
In this appendix, we derive the testable implications of the extensions to the classic Tobit and IV-Tobit discussed in Proposition 2. We discuss testable implications assuming the identification of the relevant parameters or distributions.
D.1. barros2018generalized and carson2007Tobit
barros2018generalized proposes a variant of the tobit model with elliptically contoured distributions and a non-zero threshold. At the same time carson2007Tobit proposes a Tobit model with a non-zero threshold. In this section we combine both types of results and report the testable implications with a generic non-zero threshold and a generic known or identifiable parametric distribution function. In addition, extend the latent structure to be also known or an identifiable function up to a vector of parameters but invertible in $U$. In this context, let:
@@ Appendix D Results from Proposition 2
D.2. wooldridge2005simple; honore2000estimation; honore1993orthogonality
Consider the following dynamic version of the Tobit model, which is related to wooldridge2005simple; honore2000estimation; honore1993orthogonality. Here, we specify the conditional behaviour of $c_{i}$ in the spirit of wooldridge2005simple and others.
@@
$=$ $\int[\Phi(c_{1}-\alpha_{0}-\alpha_{1}d-\alpha_{2}g(y)-c_{i})-\Phi(c_{0}-\alpha_{0}-\alpha_{1}d-\alpha_{2}g(y)-c_{i})]\phi(c_{i})dc_{i}$
@@
$=$ $\int[\Phi(-\alpha_{0}-\alpha_{1}d-\alpha_{2}g(y)-c_{i})]\phi(c_{i})dc_{i}$
Which can then be used to construct a test in a similar way as presented in the main text by adding the type of equalities $P(Y_{i,t}\geq c_{2}|D_{i,t}=d,Y_{i,t-1}=y)$, with the caveat that the right-hand side requires numerical integration or an approximation by an estimator when no closed form is available. Such estimator should ensure that the left-hand side of the equality converges at root-$N$ in order for the estimation step of the null model to be asymptotically negligible (acerenza_testing_2023, Appendix B).
@@ E.1. Non-Learnability for the classic Tobit model
As mentioned in Remark 1, the standard Tobit model is non-verifiable, that is, we cannot learn if the maintained Tobit model is the true data generating process based on the sharp equalities proposed in Section 4.1. One can show that by finding an alternative model that is compatible with the equalities (4.1)-(4.2) in all cases in which the Tobit model could not be disregarded. Thus the model is non-verifiable in the sense that we can always construct a joint probability law of $(\tilde{Y},D,\tilde{U}$) that violates the Tobit model validity but satisfies the equalities. Concretely, for the classic Tobit model suppose that,
$\left\{\begin{array}[]{lcl}\bar{Y}&=&\max(0,\bar{Y}^{*})\\ \\ \bar{Y}^{*}&=&\alpha_{0}+\alpha_{1}D+\tilde{U}\end{array}\right$. (E.4)
$\tilde{U}$ is a random variable that has the following mixed discrete-continuous “density”:
$f_{\tilde{U}|D}(\tilde{u}|d)=\begin{cases}\Phi(\hat{u})(\frac{e^{d}}{1+e^{d}})\text{ if }\tilde{u}=\hat{u}\\ \Phi(-\alpha_{0}-\alpha_{1}d)-\Phi(\hat{u})(\frac{e^{d}}{1+e^{d}})\text{ if }\tilde{u}=-\alpha_{0}-\alpha_{1}d\\ \phi(\tilde{u})\text{ if }\tilde{u}\geq-\alpha_{0}-\alpha_{1}d\\ 0,\text{ otherwise}\end{cases}$
[…]
For $c_{0}\leq Y\leq c_{1}$ for a $c_{1},c_{0}>0$ and for any value of the treatment variables $d\in\mathcal{D}$, assume that the sharp equalities hold:
@@
Now define $\bar{Y}_{d}=\alpha_{0}+\alpha_{1}d+\tilde{U}$
In the above proposed DGP, the latent model is not normal. However, the DGP is compatible with the data and the testable implications derived in Section 4 since,
$P(\bar{Y}_{d}<0)$ $=$ $P(\tilde{U}\leq-\alpha_{0}-\alpha_{1}d)$ (E.5)
$=$ $\Phi(\hat{u})\left(\frac{e^{d}}{1+e^{d}}\right)+\Phi(-\alpha_{0}-\alpha_{1}d)-\Phi(\hat{u})\left(\frac{e^{d}}{1+e^{d}}\right)$ (E.6)
$=$ $\Phi(-\alpha_{0}-\alpha_{1}d)=1-\Phi(\alpha_{0}+\alpha_{1}d)$ (E.7)
$=$ $P(Y=0|D=d)$ (E.8)
The last equality is due to the testable equalities holding. Similarly for any $c_{0},c_{1}>0$,
$P(c_{0}\leq\bar{Y}_{d}\leq c_{1})$ $=$ $\Phi(c_{1}-\alpha_{0}-\alpha_{1}d)-\Phi(c_{0}-\alpha_{0}-\alpha_{1}d)$ (E.9)
@@
Hence, when the sharp equalities hold, there exists an alternative observationally equivalent model, to the classic tobit that can induce the observed data distribution. Indeed, the observed distribution of $Y$ and the non refuted marginal distribution of $U$, do not imply that we can learn the true distribution of ($Y_{1},Y_{0},U$).